OpenStudy (anonymous):

8 1/2-2+4 3/4 A.1 3/4 B.11 1/4 C.13 1/4 D.15 3/4

OpenStudy (blackops2luvr):

i did this in 6th grade O-O

OpenStudy (anonymous):

You can change the fractions into decimals to make it easier 8 1/2=8.5 4 3/4=4.75 So now we have 8.5-2+4.75 Solve and you get 11.25

OpenStudy (anonymous):

:/.......

OpenStudy (blackops2luvr):

its b..

OpenStudy (blackops2luvr):

11 1/4 = 11.25

OpenStudy (anonymous):

she doesnt give out medals haha :/

OpenStudy (blackops2luvr):

lol

OpenStudy (blackops2luvr):

anyone guess what grade im in?

OpenStudy (anonymous):

7th

OpenStudy (blackops2luvr):

ill give a medal if u get it right

OpenStudy (blackops2luvr):

nope

OpenStudy (anonymous):

6th

OpenStudy (blackops2luvr):

nope

OpenStudy (anonymous):

9th

OpenStudy (blackops2luvr):

nope

OpenStudy (anonymous):

8th

OpenStudy (blackops2luvr):

finaly

OpenStudy (anonymous):

:(

OpenStudy (blackops2luvr):

but you dont need the medal so..

OpenStudy (anonymous):

well there goes 2 medals i didnt receive within 1 minute...

OpenStudy (anonymous):

hahaa........

OpenStudy (blackops2luvr):

lol

OpenStudy (anonymous):

stop posting on my thing i get a lot of notifications

OpenStudy (anonymous):

.

OpenStudy (anonymous):

please just stop

OpenStudy (blackops2luvr):

.

OpenStudy (anonymous):

haha ok im done.....

OpenStudy (anonymous):

How about my medal?....